### v1

# Multi-Source Urban Traffic Flow Forecasting with Drone and Loop Detector Data

Abstract
Traffic forecasting is a fundamental task in transportation research, however the scope of current research has mainly focused on a single data modality of loop detectors. Recently, the advances in Artificial Intelligence and drone technologies have made possible novel solutions for efficient, accurate and flexible aerial observations of urban traffic. As a promising traffic monitoring approach, drone-captured data can create an accurate multi-sensor mobility observatory for large-scale urban networks, when combined with existing infrastructure. Therefore, this paper investigates the problem of multi-source traffic speed prediction, simultaneously using drone and loop detector data. A simple yet effective graph-based model HiMSNet is proposed to integrate multiple data modalities and learn spatio-temporal correlations. Detailed analysis shows that predicting accurate segment-level speed is more challenging than the regional speed, especially under high-demand scenarios with heavier congestions and varying traffic dynamics. Utilizing both drone and loop detector data, the prediction accuracy can be improved compared to single-modality cases, when the sensors have lower coverages and are subject to noise. Our simulation study based on vehicle trajectories in a real urban road network has highlighted the added value of integrating drones in traffic forecasting and monitoring.
Index Terms:
Urban Traffic Analysis, Deep Learning, Multi-modal Data, Drones, Graph Neural Network.
I Introduction
Knowing the current traffic states and their future evolution is essential for traffic management and control in Intelligent Transportation Systems (ITS). Therefore, in the literature, traffic forecasting has been a primal topic of many research works. In practice, traffic monitoring and forecasting has been an important tool for transportation policy and traffic management. With accurate predictions, road users can dynamically adjust their route choices to avoid congested areas [1, 2]. City-wide traffic flow predictions can support traffic signal control [3], plannings of platoons [4], dynamic transportation system management [5] and the optimal control of autonomous buses [6]. Time-crucial services, such as emergency response, can also schedule their operations based on the predicted traffic to guarantee the arrival time [7]. These applications will improve the overall efficiency of the transportation network, reduce energy consumption, and create greater value for the economy and sustainability of society.
Owing to the limitations of data availability, many deep learning methods are built upon highway loop detector data, e.g., METR-LA and PEMS-Bay [8]. However, urban traffic forecasting presents additional challenges due to the dense and complex road network structure. Recently, Unmanned Aerial Vehicles (UAVs) or drones are emerging as a novel solution to information collection in ITS [9], as they can flexibly fly over an area of interest and record high-quality videos. At the same time, the growing power of Computer Vision, Artificial Intelligence and computational resources have made it possible to automatically and efficiently extract object information from the captured videos. These advantages have empowered drones to gather rich and accurate information from a large spatial region, making them a preferable choice for urban traffic monitoring.
Specifically, the pNEUMA experiments have started the understanding of urban congestion evolution through the analysis of video data captured by a swarm of drones over a large area in the city center of Athens [10]. The drone-captured videos present clear views of the monitored areas, and most vehicles can be identified by object detection and tracking methods [11]. Detailed trajectories collected from drones can significantly improve the understanding of various traffic phenomena, including emissions [12], multi-modal MFDs [13], path flow estimation [14], car following [15], traffic signal safety [16] and parking monitoring [17]. With vehicle trajectories, traffic variables can be accurately computed for all visible roads in the field of view (FOV) [18], resulting in a great advantage compared to grounded sensors, e.g., loop detectors and CCTV cameras, whose scopes are limited to the roads with installed facilities.
The high-quality data from drones can provide solid grounds for traffic state prediction, which will be an additional asset to the existing traffic forecasting methods based on loop detector data. However, few studies have attempted to address the traffic forecasting problem using joint observatories from drones and other information sources. Therefore, this paper aims to investigate a promising paradigm of urban traffic forecasting, where drone-measured traffic speed data are combined with loop detector data to provide high-quality speed predictions.
In summary, the main contributions are as follows. First, this paper introduces a flexible model architecture HiMSNet for road segment-level speed prediction. It consists of a local module for data fusion and a global module for spatial dependencies learning, and it is further extended with an additional branch for regional speed prediction. Second, A well-calibrated multi-source dataset, i.e., SimBarca, is created, offering traffic speed data from drones and loop detectors in the city center of Barcelona. Using the vehicle trajectories simulated in AIMSUN microsimulation software, SimBarca gathers statistics for the prediction tasks of segment-level and regional speeds. Third, Extensive experiments and analysis are provided to evaluate HiMSNet on SimBarca, providing an example for evaluating multi-source urban traffic forecasting. Finally, The methods, data and evaluations in this work are openly shared with the community to facilitate future research. (Link will be provided upon acceptance)
II Related Work
The multi-source traffic forecasting problem is an extension of the traditional traffic forecasting task. As a popular topic in transportation research, traffic forecasting has a long history and a vast collection of literature. Over the past years, the development of this field has experienced a paradigm shift from classical statistical approaches to data-driven approaches [19, 20]. Despite the great methodological differences, the critical challenge of the traffic forecasting problem remains, which is the complex temporal dynamics and spatial correlations in the data. Therefore, most deep learning methods will design specific neural network modules to tackle the dependencies in spatial and temporal dimensions.
II-A GNN-Based Methods
Since the road network has a natural graph structure, Graph Neural Networks have been extensively explored in traffic forecasting [21]. In the modern era of deep learning, the Diffusion Convolutional Recurrent Neural Network (DCRNN) [8] is one of the pioneering works to utilize a graph-based model in this area. In DCRNN, the temporal dependencies are learned by a Recurrent Neural Network (RNN), whose layer has an embedded diffusion convolution module to model the spatial dependencies by a diffusion process in the sensor graph. A network block in Spatio-temporal Graph Convolutional Neural Networks (STGCN) [22] applies a convolutional operation over the time dimension, and then propagates the information over the spatial graph structure with Graph Convolution. Following a similar design philosophy, the Temporal Graph Convolutional Network (T-GCN) [23] propagates neighborhood information over the graph at each time step and then applies an RNN to learn the temporal dependencies.
However, the correlations in the data are often beyond the description of a pre-defined graph structure, which is typically an adjacency matrix based on geographical distance or road topology. Therefore, many subsequent works have focused on learning an adaptive graph structure from the data, instead of following a fixed graph. Graph WaveNet [24] develops an adaptive graph to learn the spatial relations between the sensors, which is achieved by composing the graph with learnable positional embeddings. The Attention-based Spatial-Temporal Graph Convolutional Networks (ASTGCN) [25] further introduces an attention-based mechanism to both the spatial and temporal dimensions, which effectively assigns different weights to the neighbors according to the correlation patterns. The Adaptive Graph Convolutional Recurrent Network (AGCRN) [26] also employs an embedding-based adaptive graph as Graph WaveNet [24], but proposes to learn a normalized adjacency matrix instead of the adjacency itself.
This spatio-temporal framework has been further developed in the literature. In Multi-variate Time Series Graph Neural Network (MTGNN) [27], a node will exchange information with neighborhoods at different distances, using the Mixhop [28] technique. The Graph Multi-Attention Network (GMAN) [29] has an encoder-decoder structure equipped with an attention mechanism for space and time dimensions, and the encoder message is further enhanced by cross attention between historical and future time step embeddings. TwoResNet [30] addresses regional traffic with a low-resolution module at macroscopic scale, and the microscopic traffic is handled by a high-resolution module. The Traffic Graph Convolution (TGC) network [31] adopted intuition from transportation and defined the adjacency matrix based on the travel time between the spatial locations, instead of geographical distance. The Multi-Weight Traffic Graph Convolution (MW-TGC) network [32] integrated various metadata into the design of adjacency matrix, including geographical distance, angle, speed limit and number of paths between locations.
II-B Transformer-Based Methods
In recent years, the Transformer [33] model has also been introduced into traffic forecasting. The core mechanism of Transformer is attention, which computes pair-wise similarities between two sets of elements and then aggregates the information based on the similarities. Such a generic design makes Transformers suitable for the adaptive learning of both spatial and temporal dependencies in traffic forecasting problem. Thus, this has become a shared foundation for many transformer-based methods.
Yan et al. [34] use a transformer model to learn the spatial dependencies, and employ an LSTM [35] for temporal learning. The adjacency matrix is used in transformer decoder part to enforce the spatial constraints. Xu et al. [36] propose to learn temporal relations with transformers, and fuse the output of transformer and Graph Neural Network (GNN) to obtain a joint representation of the spatial relations. In this way, the model can make use of the flexibility of pair-wise correlations in the transformer model, and also follow the spatial dependencies of the graph structure. Similar to GMAN [29], the Meta Graph Transformer (MGT) [37] designs a cross-attention between encoder and decoder, which allows the model to capture the relations between historical and future time steps. To enhance the graph structure, MGT also encoded multiple weight matrices based on geographical distance, historical traffic flow and vehicle counts. Shao et al. [38] adapted the method from Masked Auto-Encoder [39], and propose to obtain a better data representation by a pretraining phase on multiple datasets, where a transformer model is trained to complete the missing values with only partial inputs. PDFormer [40] clusters segments of traffic time series into sets of representative patterns, and utilize them to identify similar patterns in new input sequences when making predictions.
Since traffic forecasting datasets usually do not have a big scale, building a complicated model is not always necessary. For example, STID [41] achieves similar accuracy as PDFormer by using MLPs only. A later study on general time series forecasting has also shown that linear models can achieve similar performance as transformer models [42] if properly tuned. Compared to existing literature, this study focuses on an urban area with denser road networks and highly dynamic traffic. At the same time, multiple data sources are utilized together to provide more accurate and robust predictions. In previous studies, the research scope is often the traffic flow prediction on highways using data from loop detectors only. This work aims to provide a simple yet effective model along with comprehensive evaluations to lay the foundation for future research. Therefore, we refrain from using a complex model design and keep the implementation concise.
III Methodology
This section will first introduce the formulation of multi-source traffic forecasting. Then a baseline model for multi-source traffic forecasting will be presented, and the final part will describe the dataset creation method using vehicle trajectories.
III-A Multi-Source Traffic Forecasting
In a modern transportation network, there are various types of sensors (information sources) providing traffic information at different locations in different ways. If the network traffic states $\mathbf{X}$ are observed at time $t$, location $p$ using method $m$, and the measured value is $v$, this measurement procedure can be described as
$v=\mathcal{O}_{m}(\mathbf{X},t,p),$ (1)
where $\mathcal{O}_{m}$ represents the general physical or statistical process to obtain the measurement.
For a real-world urban transportation network, the accurate and true state $\mathbf{X}$ is often unknown or difficult to determine. The practically available information is what the sensors provide, for example, one such measurement can be noted as $(t,v,m,p)$. Therefore, the observations collected over time at location $p$ using method $m$ can be noted as a time series:
$\mathcal{S}^{m,p}=\left\{(t_{0},v_{0}),(t_{1},v_{1}),...,(t_{n},v_{n})~{}|~{}m,p\right\},$ (2)
where the time steps $t_{i}$ are chronologically sorted. Such a time series has its own evolution patterns over time, and at the same time is correlated with time series from other locations, since they belong to the same transportation network $\mathcal{G}$. Besides, since different observation methods reveal different aspects of the traffic states, i.e., different sensor coverages, the inter-modality correlation is also important in multi-source traffic forecasting. As a result, if there are $M$ observation methods and $P$ locations in total, a Multi-Source Time Series (MSTS) is the collection of time series from all modalities and locations, which comprehensively describe the available traffic information in the transportation network.
$\mathcal{S}=\left\{\mathcal{S}^{i,j}\right\},\text{where}~{}i\in[1,2,...,M],~{}j\in[1,2,...,P].$ (3)
An MSTS is considered synchronized if, for any time step with records, the measurements with all modalities on the same variable are available at all locations. This will be the special case where all the road segments of interest are covered by all types of sensors, and these sensors are designed to provide measurements at the same time steps. However, this is not realistic mainly because different sensors can have different update strategies and spatial coverages. As a result, the MSTS for a transportation network is usually unsynchronized, and it cannot be simply concatenated to be a time series with multiple variables. Instead, the spatial correlations and modality characteristics should be considered in the model.
With this concept, a dataset $\mathcal{D}$ is a collection of MSTS $\mathcal{S}_{k}$, the corresponding labels $\mathbf{y}_{k}$ and the road graph $\mathcal{G}$, with $k$ being the index of the sample.
$\mathcal{D}=\{(\mathcal{S}_{1},\mathbf{y}_{1}),(\mathcal{S}_{2},\mathbf{y}_{2}),...,(\mathcal{S}_{N},\mathbf{y}_{N});\mathcal{G}\}.$ (4)
Suppose a neural network $f$ parameterized by $\theta$ is trained to predict $\mathbf{y}$ using the input MTMS $\mathcal{S}$ and the road graph $\mathcal{G}$, the training loss can be generally noted as:
$L(\theta,\mathcal{D})=\mathbb{E}_{(\mathcal{S},\mathbf{y})\in\mathcal{D}}\left[l(\mathbf{y};f_{\theta}(\mathcal{S},\mathcal{G}))\right],$ (5)
where $l$ is a distance function to measure the differences between predictions and labels. The exact form of the label $\mathbf{y}_{i}$ and loss $l$ will depend on the task (e.g., forecasting, interpolation and imputation), and more details will be provided in Section IV-B, where the experiment details will be elaborated. Besides, the notations of time series sequences and dataset are inspired by [43].
III-B Model Architecture
Figure 1 shows the structure of our baseline model Hierarchical Multi-Source Network (HiMSNet) for traffic forecasting. This architecture consists of a Global Message Exchange (GME) module to learn the spatial correlations between different locations, and a Local State Evolution (LSE) module to handle temporal dependencies and multiple data modalities at each location. It is a simple and straightforward baseline model that addresses the spatial, temporal and modality correlations in multi-source traffic forecasting.
Following the road network, the input data can be naturally organized as a graph structure, where a node corresponds to a road segment, and an edge indicates the connection between two road segments. In the LSE module, the temporal encoder $\mathcal{H}_{e}$ will take $\mathcal{S}^{*,p}$, i.e., the time series from all observation methods at location $p$ (modalities marked in colors), and separately encode the time series of each modality into a unified representation $\mathbf{z}_{p}$. Then, $\mathbf{z}_{p}$ will go through a message encoding layer $\mathcal{R}_{1}$ and become a message to be shared with other locations. The GME module will then take the messages from all locations ($\mathbf{Z}$), use its encoder $\mathcal{F}_{e}$ and decoder $\mathcal{F}_{d}$ to compute the new feature with exchanged information ($\mathbf{Q}$). The message decoding layer $\mathcal{R}_{2}$ will take the element in $\mathbf{Q}$ corresponding to location $p$, and concatenate it together with the elements in $\mathbf{z}_{p}$ to form a joint feature $\mathbf{q}_{p}$, which contains the information from the location $p$ itself and its vicinity. Finally, the temporal decoder $\mathcal{H}_{d}$ will take $\mathbf{q}_{p}$ and predict the future traffic states $\hat{\mathbf{y}}_{p}$.
Figure 1: The model structure of HiMSNet. The model consists of a global layer and a local layer. The local layer processes the time series from the sensors that are monitoring a road segment, and obtains a joint feature for each road segment. The global layer will initiate message exchanges across the road network, and enhance the local joint features with vicinity information.
Figure 2: Baseline branch for regional prediction. This branch gathers the features of all road segments in a region and predicts the regional traffic.
HiMSNet is flexible and can be easily extended with an additional branch (Figure 2) for regional traffic forecasting, which can be helpful in regional traffic control. Given a defined region $I$, the regional branch will take the average of joint feature $\mathbf{q}_{p}$ for all $p\in I$ as the regional representation $\mathbf{q}_{I}$, and predict the regional average traffic state $\hat{\mathbf{y}}_{I}$ with the prediction module $\mathcal{H}_{I}$.
III-C Dataset Creation
Although existing drone datasets have already provided high-quality vehicle trajectory data, e.g., pNEUMA [10] and highD [44], the aim of this work is to predict the traffic states for all road segments within a large urban region, which is beyond the scope of available datasets. For this reason, detailed vehicle trajectory data are generated using the microscopic traffic simulator Aimsun and its road network for central Barcelona [45]. The simulation environment has been properly calibrated based on real-world traffic scenarios, and it has been applied in various analyses on urban traffic, such as network wide-perimeter control [46, 47]. As shown in Figure 3, the network consists of many road segments, intersections as well as centroids ($\odot$ like) that serve as origins and destinations of vehicles. The fine details from microscopic traffic simulations allow great flexibility for subsequent processing and analysis, e.g., the computation of vehicle flow, density and speed. Concretely, there are mainly three types of available information:
  • 1.
    Network topology and the geometry of road segments.
  • 2.
    Vehicle locations at each simulation time step, referenced within road segments.
  • 3.
    The time when vehicles enter or exit a road segment.
Figure 3: The simulated urban transportation network with detailed views of an intersection and a road segment (ID 9971), which will be used as visualization example in the experiments.
In the simulation runs, the vehicles are created according to a specific OD matrix calibrated using historical data. It is a square matrix with elements representing the number of vehicles going from an origin centroid to a destination in the time period of the demand. However, the simulator provides only one such matrix, which can limit the diversity of the generated trajectory data and the ability of the machine learning model. Therefore, the non-zero elements in the OD matrix are randomly augmented by the following operations, while the zero elements are kept unchanged:
  • 1.
    randomly set an element to zero
  • 2.
    increase or decrease an element by a random percentage
  • 3.
    multiply all elements by a common random demand scale
When computing traffic variables from trajectories, the characteristics of different sensors need to be carefully considered to resemble the real-world data collection process. In this work, we focus on the speed measurements from two types of sensors: inductive loop detectors and drones. Inductive loop detectors can only measure vehicle speeds at certain points, as they are installed at fixed locations. Since vehicles can change their speeds within a road segment, the speed measurements of a loop detector are valid only in its proximity and often cannot represent the situation of the whole segment. On the contrary, drones can capture nearly all vehicles in their field of view, therefore it is possible to compute accurate segment-level speeds by using the complete vehicle trajectories. In this paper, we refer to these two types of speeds as point speed and segment speed respectively, and both of them can be extracted with the trajectory data from microscopic simulation.
Figure 4: Traffic variable extraction with trajectory splits
Figure 4 shows an example of a road segment along with the trajectory of one vehicle. At each simulation time step ($t_{1}\sim t_{4}$), the vehicle position and speed are known. However, this work intends to develop a preprocessing pipeline also applicable to real-world drone datasets, and thus will avoid using the speed information and assume constant speed between the two known positions. Because in reality, the vehicle speeds are not directly measured from drone videos, but are calculated using the vehicle locations at different frames. Still, for the entry and exit points, only the timestamps ($t_{in}$ and $t_{out}$) are known, thus the positions have to be extrapolated from the closest known positions (e.g., the vehicle moves at speed $v_{1}$ from $t_{in}$ to $t_{1}$).
Then, two adjacent points can compose a trajectory split $(i,t_{s},t_{e},x_{s},x_{e})$, which means vehicle $i$ travels from position $x_{s}$ to $x_{e}$ between time $t_{s}$ and $t_{e}$. Afterward, the trajectory splits can be grouped by different time resolutions (e.g., every 5 s for drones and 3 minutes for loop detectors) to simulate the update frequency of drone data and loop detector data, or by spatial region to provide regional traffic data. For segment speed, we sum the travel distance $\Delta x=x_{e}-x_{s}$ and travel time $\Delta t=t_{e}-t_{s}$ of all trajectory splits in the road segment, and compute the segment speed following Edie’s definitions of generalized flows and densities [48]:
$\bar{v}_{s}=\frac{\sum\Delta x}{\sum\Delta t}.$ (6)
For a loop detector installed $d$ meters from the start, if a trajectory split has $x_{s}<d<x_{e}$, then the corresponding vehicle is detected at the timestep of the trajectory split. We count the number of detected vehicles $n$ in a time interval (e.g., every 3 minutes), and arithmetically average their speeds to get point speed as
$\bar{v}_{p}=(\sum\frac{\Delta x}{\Delta t})/n.$ (7)
By choosing different time resolutions for Equation 6 and 7, the simulated loop detector and drone measurements can be computed individually for each road segment. Similarly, regional speed can be obtained by applying Equation 6 to all the trajectory splits in a region during a time interval.
Then, the training samples for the prediction model can be generated using a sliding window approach as depicted in Figure 5. Each simulation session lasts for 4 hours in the simulation world, with a 15-minute warm-up and 1 hour 45 min main demand, while the real-world run time differs according to the traffic demand. This demand profile is intended to simulate the morning peak hours in the road network. The sliding window will start at the end of the warm-up period up to 2 hours after the beginning of the main demand, and the window size is 1 hour with a 3-minute step size between consecutive samples. Inside a window, the first half is used as input and the second half is used as targeted output for the prediction model. Therefore each simulation session will have 20 training samples, each with 30-minute input and 30-minute output.
Figure 5: Training sample extraction using a sliding window approach
IV Experiments
This section will present the dataset description at both micro and macro levels, implementation details, and the evaluation of prediction performance.
Generally, SimBarca introduces two input modalities, i.e., vehicle speeds from loop detectors and drone monitoring, and two prediction tasks, i.e., road segment speed prediction and regional speed prediction. As depicted in Figure 5, the prediction model is required to forecast the future 30 minutes given the data of the past 30 minutes. The input modalities provide measurements of road segments with different time resolutions, where loop detectors report speeds every 3 minutes and drones report speeds every 5 seconds. Both prediction tasks are labeled with future speed values every 3 minutes. The segment speeds computed by Equation 7 are used as segment-level labels, and the regional labels are obtained by aggregating all trajectory splits within a region and applying the same formula. When computing the point speed, the loop detector is assumed to be installed at the midpoint of a road segment.
Although in reality, drones are able to provide high-resolution video at 30 frames per second, the high data volume will create strong needs for data transfer and computation in a real-time framework. Therefore, this work simplifies the process with a lower temporal resolution that still produces accurate speed estimations from a traffic engineering perspective.
In total, SimBarca consists of 101 simulation runs, including one with the original demand profile unmodified and the other 100 with different demands augmented from the original one. Out of all the simulation sessions, 75 runs are randomly selected for training and the rest are for testing. In the graph-structured data, there are 1570 nodes and 2803 edges in total, and the adjacency matrix is a symmetric binary matrix by assuming the graph is undirected. The experiments start with a full-information setting where the model has access to segment speeds and point speeds for all road segments (Section IV-C and IV-E), and then move on to a more realistic scenario where the model only has access to noisy data with partial spatial coverage (Section IV-F).
IV-A Data Visualization
In the first place, the traffic statistics required by Equation 6 and 7, e.g., total distance and time traveled by all vehicles, are calculated at each simulation time step (0.5s), and then utilized for speed computation. Figure 6 compares the segment speed and point speed per simulation time step, which are the finest possible details from the simulation data. The segment speed is computed with all the vehicle trajectories in the road segment, and therefore it can reflect the overall traffic condition in the segment. There is a strong periodic pattern in segment speed due to traffic lights, with which one can identify the acceleration and deceleration patterns in the segment speed. On the other hand, the point speed can only capture the instant speed of vehicles at a specific location, and thus contains more stochastics due to randomized vehicle passing. As a result, from the aspect of congestion development and propagation, point speed does not provide as much information as the segment speed. Besides, the randomized simulations have resulted in varying dynamics in the transportation network. For example, Figure 7 shows the segment speed of the same road segment from multiple simulation runs. The periodic pattern is consistent across different runs, but the details vary because of the demand augmentation and vehicle generation processes in the simulation.
Figure 6: Comparison of segment speed and point speed per simulation time step (0.5s), segment ID 9971
Figure 7: segment speed from multiple simulation runs, segment ID 9971
The initial statistics are then aggregated to coarser time resolutions for the training of the prediction model. In all the experiments, the model is trained to forecast the future segment speed rather than future point speed, as the former can better reflect the traffic dynamics in the whole road segment. The differences between point speed and segment speed are still clear in the aggregated data. For example, Figure 8a shows the segment speed and point speed for the road segment with ID 9971 (as marked in Figure 3), both aggregated every 3 minutes. In this segment, point speed overestimates true segment speed roughly by 5 m/s.
Figure 8b shows an example training sample obtained by the sliding window approach. In the input sequences (first 30 minutes), the segment speed (Drone Input) is given every 5 seconds, where the periodic pattern can be visually identified as the per-time-step case in Figure 6. For example, the segment speed drops to zero periodically at every cycle of traffic signals due to the effect of the red phase. The point speed (LD Input) is aggregated to a 3-minute interval as common practice, and still it overestimates the true speed, since the measured loop detector speed is influenced by the distance of the sensor from the stop line. The second 30 minutes of the figure shows the prediction labels, including the segment-level future speed (Segment Label) as well as the future speed of the region (Regional Label) that this segment belongs to. In this example, traffic congestion has already built up within the short time span of the training sample, as both the segment and regional speed have decreased.
Figure 8: Visualization of aggregated time series (a) and training sample (b) for road segment with ID 9971
In the regional prediction task, the road segments are clustered into 4 regions based on spatial distances using K-Means for simplicity. More advanced clustering algorithms could also be utilized for this task to consider both road structure and historical traffic data [49]. In Figure 9, the center point of each road segment is marked with a point, and the color represents the region it belongs to. The map is also discretized into grids with cell size $220~{}\times~{}220m$, and it is assumed that a drone hovering at the center of a grid can monitor the traffic condition of all road segments, whose center points are inside the grid.
Figure 9: Regions for regional prediction and grids for drone monitoring
Figure 10 shows the Macroscopic Fundamental Diagram (MFD) of the whole road network under different demand scales, which is the augmentation factor multiplied with the demand matrix as mentioned in Section III-C. Notably, the exact number of vehicles in the simulation is still affected by the other two augmentation methods, and thus is not strictly proportional to the demand scale. In these plots, the vehicle flow is computed by $q=D/tL$, where $D$ is the total distance traveled by all vehicles during the time interval of concern $t$, and $L$ is the total length of the road network. Similarly, the vehicle density is obtained using $k=T/tL$ with $T$ being the total travel time of all vehicles. Each subplot shows the MFD with $t$ being 5 seconds and 3 minutes, corresponding to the time resolution of drone and loop detector data respectively, and the data points are annotated with the time from the simulation start.
Figure 10: MFD under different demand scales, (a) 120%, (b) 150% and (c) 180%.
These plots show that as the demand level increases, the vehicle flow becomes saturated earlier and the vehicle density reaches a higher maximum value. Meanwhile, the network-level average speed also decreases greatly as the congestion becomes more severe. Besides, the hysteresis loop is consistent with previous empirical and simulation observations [50]. The increased congestion can be further evidenced by the travel time distribution of the network, as shown in Figure 11. The increase in demand scale leads to a larger average and a longer tail in the travel time distribution, which means more and more vehicles are experiencing longer travel times.
Figure 11: Network travel time distribution under different demand scales, (a) 120%, (b) 150% and (c) 180%
IV-B Implementation Details
To provide a reasonable baseline model, the implementation of HiMSNet is kept as concise as possible. The first step in HiMSNet is value embedding, i.e., the values and corresponding timestamps are embedded into a feature space using a linear layer, whose output dimension is 64. Specifically, the missing values, which appear when a road is not monitored or no vehicle passes a monitored road, are replaced with learnable neural network parameters of equal size. After the embedding, the drone modality has 2 layers of 1D convolution with kernel size 3 and stride 3 to down-sample the temporal resolution. Then, each modality has an individual temporal encoder $\mathcal{H}_{e}$ consisting of 3-layer LSTMs with hidden size 64, and their last step outputs will be used as the local feature $\mathbf{z}_{p}$. The decoders for both segment-level and regional predictions ($\mathcal{H}_{d}$ and $\mathcal{H}_{I}$) are 2-layer MLPs with doubled hidden size (128) and output dimension 10 to predict 10 future time steps (30 minutes). Contrary to modality-specific temporal encoders, the decoder is shared as the features have been combined. The message encoding layer $\mathcal{R}_{1}$ is a 1 $\times$ 1 convolution layer that projects the features while keeping the dimension unchanged. Accordingly, the message decoding layer $\mathcal{R}_{2}$ inherits the same structure. The GME module is implemented with 3 layers of GCNConv [51] followed by ReLU activation and Layer Normalization [52].
In the experiments, Mean Absolute Error (MAE) is used as the loss function for both segment-level and regional prediction tasks, and equal weights are assigned to the two tasks by default. HiMSNet is trained for 30 epochs using Adam optimizer with a learning rate 0.001, weight decay $1\times 10^{-4}$, $\beta=(0.9,0.999)$ and batch size 8. We use a learning rate scheduler that gradually increases the learning rate from 0 to the chosen value during the first epoch, and decreases it to 10% and 1% of the chosen value at 70% and 85% of the total epochs. All implementations are based on PyTorch, and the experiments are conducted on a computer equipped with an Nvidia RTX 4090 GPU.
IV-C Overall Prediction Performance
Following the common practice in traffic forecasting literature [8], the prediction results are evaluated with three metrics: MAE, Root Mean Square Error (RMSE) and Mean Absolute Percentage Error (MAPE). Their vector forms for a pair of prediction ($\hat{\mathbf{y}}$) and ground truth ($\mathbf{y}$) are defined as follows:
$\begin{cases}\text{MAE}&=\frac{1}{n}\sum_{i}^{n}|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}|\\ \text{RMSE}&=\sqrt{\frac{1}{n}\sum_{i}^{n}(\hat{\mathbf{y}}_{i}-\mathbf{y}_{i})^{2}}\\ \text{MAPE}&=\frac{1}{n}\sum_{i}^{n}|(\hat{\mathbf{y}}_{i}-\mathbf{y}_{i})/\mathbf{y}_{i}|\end{cases},$ (8)
where $n$ is an index for flattened predictions and labels.
In the evaluation, the speed values take the unit of m/s, and the metrics are calculated for each road segment (or region) and then averaged over all segments (regions). Since a zero segment speed in the label will result in infinite MAPE, we only evaluate MAPE when the speed value is greater than 1 m/s, and we refer to this modified metric as MAPE*. These metrics are reported for 15 min and 30 min prediction windows for both segment-level and regional task, which correspond to time steps 5 and 10 in the dataset.
Table I compares the prediction errors of HiMSNet with three simple baselines, best results in bold. Last Observation (LO) uses the last observed speed as the segment-level prediction, while Input Average (IA) uses the average speed of the input sequence. Both of them can be configured to use loop detector (ld) or drone (drone) data, and the regional prediction is made by averaging the segment-level predictions. The Label Average (LA) baseline uses the average speed of the whole test set labels as prediction, which applies to both prediction tasks. In fact, the LA baseline is the best possible constant-value prediction that leads to the lowest MAE. The second group of methods are variants of HiMSNet, where we present a brief ablation study to examine the importance of different components. The exact settings are: disable the GME module (/gnn) to exclude vicinity information, only use loop detector modality in the input (ld), and only use the drone modality (drone) in the input.
The IA and LO baseline with loop detector data have significantly higher errors than the variants using drone data, which originates from the difference between loop detector measurements and the drone measurements as shown in Figure 6. IA(drone) also outperforms LO(drone), which is expected as the drone-measured speed has a high update frequency (every 5s) to reveal the traffic dynamics within a signal cycle. Therefore using the last observation as a prediction will be very biased towards the end of the cycle. The LA baseline has the lowest error among the first group of results, and HiMSNet outperforms all the baselines, especially in the regional prediction task.
TABLE I: Prediction errors for road segment task and regional task
Model Segment 15 min Segment 30 min Regional 15 min Regional 30 min
MAE MAPE* RMSE MAE MAPE* RMSE MAE MAPE* RMSE MAE MAPE* RMSE
IA(ld) 4.50 142.7% 5.33 4.58 144.2% 5.45 6.02 174.7% 6.07 6.18 192.8% 6.27
IA(drone) 1.53 42.3% 2.06 1.70 45.7% 2.32 2.38 72.4% 2.51 2.54 83.9% 2.75
LO(ld) 4.47 141.6% 5.32 4.55 143.3% 5.44 5.95 172.2% 5.99 6.11 190.5% 6.02
LO(drone) 3.26 90.5% 4.23 3.33 92.4% 4.34 2.70 80.4% 2.78 2.87 92.3% 3.01
LA 1.24 28.4% 1.90 1.43 31.7% 2.18 0.83 26.8% 1.08 1.01 34.3% 1.34
HiMSNet 1.01 23.2% 1.68 1.17 26.3% 1.94 0.20 6.1% 0.28 0.30 9.9% 0.46
HiMSNet/gnn 1.03 24.0% 1.74 1.21 27.4% 2.03 0.20 6.2% 0.28 0.31 10.5% 0.47
HiMSNet(drone) 1.00 23.2% 1.68 1.16 26.3% 1.93 0.19 6.0% 0.27 0.29 9.6% 0.45
HiMSNet(ld) 1.13 26.3% 1.84 1.29 29.3% 2.10 0.55 15.5% 0.63 0.75 22.2% 0.88
For the regional task, most of the HiMSNet variants have close prediction performance, except that HiMSNet(ld) clearly has higher errors than the other variants. Moreover, although the drone modality has been excluded from the input of HiMSNet(ld), the training labels of this model are still computed using the vehicle trajectory data, and in reality, such data are not possible with loop detectors only. Compared to using drone data directly as input, HiMSNet(ld) faces additional challenges because of the difference between point speeds and segment speed (Figure 6). This performance difference suggests that the translation from point speed to segment speed can not be easily learned, and also highlights the necessity of drone data modality.
Similarly, in the segment-level prediction task, the model using only drone inputs HiMSNet(drone) has the lowest errors among all the HiMSNet variants, and HiMSNet(ld) has the highest. This observation again highlights the importance of the drone modality, and also verifies the difficulty of learning the translation from point speed to segment speed. Under the full-information setting, the single-modality HiMSNet(drone) even slightly outperforms the multi-modality HiMSNet, suggesting that the loop detector modality may be uninformative when the same locations are covered by drones. Different from the regional task, the inclusion of vicinity features shows higher impacts on the segment-level task, as removing the GME module in HiMSNet(/gnn) results in higher prediction errors, compared to the default HiMSNet.
IV-D Detailed Analysis of Prediction Performance
Figure 12: 30-min-ahead prediction for a road segment (ID 9971) at different simulation sessions. In each subplot, the $x$-axis is the time step index (3 min per step), and $y$-axis is the speed value in m/s. The title indicates the number of the simulation session (Sim.) out of the total 101 sessions, along with the demand scale (D.S.) of the session.
Traffic management and control often require predictions ahead of time to allow timely actions, and therefore the 30-min-ahead prediction is of particular interest. Figure 12 shows the 30-min-ahead prediction of HiMSNet for one road segment with the groundtruth, where each subplot corresponds to a different simulation session. As shown in the plot, the groundtruth can often have rapid changes over time, which reflects the dynamic nature of urban traffic flow. Due to different traffic demands, the speed of a segment also changes vastly from one simulation run to another. For example, in Session 7 of Figure 12, the segment speed is relatively stable (mostly around 5 m/s), but in Session 86, the speed decreases from 2.5 down to 0 because of heavy congestion. Since the time step is 3 minutes, the prediction horizon of 30 minutes corresponds to 10 steps. The results show that HiMSNet can capture the general trend of the speed changes (Session 11), but at the same time, it has difficulties predicting the rapid changes in the segment speed. The model can predict congestion well in advance (Session 53), but there are also cases where HiMSNet only knows the congestion after it happens (Session 73).
Figure 13: Average test set MAE (unit m/s) for different locations, coordinate format EPSG:32601
Figure 13 illustrates the spatial distribution of the prediction MAE of HiMSNet on SimBarca test set. This MAE is averaged over all the prediction horizons, which indicates the overall performance. Most of the road segments have low prediction MAE, but there are still a few places with higher errors, which suggests the traffic states there are more difficult to predict. From the perspective of traffic management, these locations are probably worth close monitoring.
To further analyze the error distribution and identify the challenging scenarios in urban traffic forecasting, Figure 14 shows the box plots for the average MAE over all prediction horizons grouped by different factors. First of all, the prediction error increases as the demand scale increases, and a scale at 180% will lead to significantly higher errors. This trend is evident in both segment-level (Figure 14a) and regional tasks (Figure 14b). This is consistent with the intuition that a higher traffic demand will lead to heavier congestion and thus will induce more dynamics in the traffic flow, making the prediction more challenging. Figure 14c shows that the road segments with intermediate average speeds are harder to predict, while the ones with very low or very high average speeds have much lower errors. This is probably because predicting the speed of a road segment will be easy when the road segment is always congested or always free-flowing. However, when it often has a changing status (and thus intermediate average speed), accurate speed predictions will require the model to punctually capture the dynamics in the network traffic flow, which is more challenging.
Figure 14: Average MAE (unit m/s) over all prediction horizons grouped by different factors. (a) by demand scale for the segment-level task, (b) by demand scale for the regional task, (c) by average segment speed over the test set (segment-level task).
IV-E Effects of Hyperparameters
The choices of hyperparameters can affect the performance of the model. Figure 15 compares the 30-min-ahead prediction MAE under different hyperparameter settings, including adjacency hops, loss balance between tasks, training epochs and model sizes. The ablation study in Table II has shown that vicinity information is important for the segment-level prediction task, which highlights the effectiveness of the GME module. Still, the scope of the considered neighborhood should be properly chosen. In Figure 15a, a single GCNConv layer with 1-hop adjacency will strictly follow the road network definition and only take into account the immediate neighbors of a road segment. Although this message passing scheme is stacked 3 times in the GME module, it still turns out to be suboptimal due to the limited adjacency scope. On the contrary, a layer with 5-hop adjacency will have a large scope and include the nodes that are 5 hops away from a road segment, and thus it may introduce much irrelevant information. Therefore in the three experimented setups, the 3-hop adjacency layer achieves the best performance for the segment-level task, as it has the most suitable vicinity scope.
Figure 15b presents the regional and segment-level prediction errors when the weight of the regional task is set to different values, and the weight of the segment-level task is always fixed at 1. When the regional task has a low loss weight, the model will be able to focus more on the segment-level task and thus achieve lower errors there. However, the trend in the regional task is unclear when it is either underweighted or overweighted, and therefore in all other experiments, both tasks are chosen to have equal weight of 1 for simplicity. Figure 15c compares the prediction errors at different training epochs, and a longer training duration will generally lead to lower error for the segment-level task but the regional task will be compromised. Therefore in all other experiments, a duration with 30 epochs is chosen for both simplicity and a balance of the two tasks. Figure 15d shows the prediction errors when the model size varies, where a larger model size can not lead to better performance. For this reason, a default model size of 64 has been chosen for the model.
Figure 15: 30-min-ahead prediction MAE (unit m/s) under different experiment settings. (a) by adjacency hop, (b) by loss weights, (c) by the number of training epochs and (d) by the hidden dimension.
IV-F Partial Observation and Noisy Data
In previous experiments, the model has access to the speed data of all road segments, and all data are free from noise. However, a city can not install loop detectors on every road segment, or fly drones to cover the whole city all the time, and measurements from these sensors are often subject to noise. Therefore, this section investigates a more realistic scenario with partial observations and noisy data, abbreviated as PN.
For partial observations, the loop detector modality has 10% coverage, i.e., 157 road segments out of the 1570 segments are equipped with loop detectors. The segments with loop detectors are randomly selected after filtering out the locations with more than 10% invalid values. The same set of loop detector locations are applied to all simulation runs (both training and testing), since they are installed infrastructure and can not be changed easily. As for drones, it is assumed that each drone can cover a grid cell of $220~{}\times~{}220m$ while following common regulations on flying height, and 21 drones (also 10%) are randomly placed over the 212 non-empty cells in Figure 9. Contrary to loop detectors, drones are flexible monitoring agents, and they are randomly relocated every 3 minutes, without considering the time to move between cells.
Multiplicative Gaussian noise is added to the segment-level speed data. The noises have zero mean and standard deviation 0.05 and 0.15 for loop detector and drone modality. Notably, partial observation and noisy data are introduced in the training and evaluation phase in different ways. In the training phase, both the inputs and segment-level labels are partial and noisy. Meanwhile, the regional labels are computed with partial segment-level details, i.e., total vehicle travel distance and time in observed road segments. During evaluation, the input data remain partial and noisy, but the labels are fully visible and noise-free. This approach is intended to assess the model’s ability to infer the traffic dynamics of the entire road network even with access to only partial and noisy data.
In all the experiments, HiMSNet is trained to predict the future segment speed and regional speed, which are not available from loop detectors. Therefore, the models HiMSNet(ld) and HiMSNet_PN(ld) have also benefited from the drone data in the training phase, despite that their inputs have loop detector data only. In fact, they are trained to predict future drone-measured traffic states by using past loop detector data only. This corresponds to a scenario where the training data are collected using drones, but no drones will be available after the prediction model is deployed. As a result, it is more challenging than using drone modality directly in the input, because the models need to cope with the different evolution patterns of the two data sources.
In reality, it is also possible when drone data are available neither in training nor testing, and the training labels are obtained from trajectories of taxis or ride-hailing vehicles. Although it is still possible to obtain segment speed and regional speed labels, the data quality will be much lower compared to drones, due to low penetration rates. This scenario is denoted by HiMSNet_PN(ld-), as the drone modality is considered to be absent in both training and testing phases. For implementation, the noise in training labels has zero mean and doubled standard deviation at 0.3.
TABLE II: Prediction errors with partial observation (P) and noisy (N) data
Settings Segment 15 min Segment 30 min Regional 15 min Regional 30 min
MAE MAPE* RMSE MAE MAPE* RMSE MAE MAPE* RMSE MAE MAPE* RMSE
HiMSNet_PN 1.06 24.5% 1.75 1.21 27.4% 1.99 0.32 10.0% 0.43 0.44 15.1% 0.63
HiMSNet_PN/gnn 1.14 26.2% 1.90 1.32 29.3% 2.18 0.34 10.6% 0.47 0.47 16.1% 0.67
HiMSNet_PN(drone) 1.07 24.7% 1.77 1.22 27.6% 2.01 0.32 10.1% 0.44 0.44 15.4% 0.64
HiMSNet_PN(ld) 1.12 26.0% 1.83 1.28 28.8% 2.07 0.41 13.3% 0.57 0.52 18.6% 0.76
HiMSNet_PN(ld-) 1.16 25.8% 1.83 1.32 28.6% 2.07 0.56 18.6% 0.86 0.72 25.0% 1.07
Table II presents the prediction errors of HiMSNet under the partial observation and noisy data settings, best results in bold. When compared with their full-information counterparts in Table I, the PN group generally has higher prediction errors, which is expected as the model has less information and more noises to deal with.
Among all the PN settings, the model with both data modalities HiMSNet_PN has the best performance, followed by the drone-only model HiMSNet_PN(drone) and loop detector-only model HiMSNet_PN(ld). This result suggests that the drone data is a better source for urban traffic forecasting, and the loop detector data can still provide complementary information when the drone data is only partially available. However, under the full-information setting, the loop detector data shows a slightly negative effect as the single-modality setting with drone data outperforms the two-modality default setting on the segment-level task. Besides, the model with partial and noisy loop detector data HiMSNet_PN(ld) does better than its counterpart with full loop detector data HiMSNet(ld). Still, in the more realistic partial data scenario, when both modalities are present, their coverages can be complementary, and the loop detector data can provide useful information to the model and improve the prediction performance. The importance of spatial correlation manifests when the data is partial and noisy. In the segment-level task, the accuracy difference between HiMSNet_PN and HiMSNet_PN/gnn is much larger than the difference between HiMSNet and HiMSNet/gnn, which highlights the necessity of the GME module in the model. The necessity of drone modality can be further emphasized by comparing HiMSNet_PN(ld-) with HiMSNet_PN(ld), as the prediction errors have increased significantly for the regional task when the drone data is totally absent.
Figure 16 shows the 30-min-ahead prediction of HiMSNet_PN for the road segment with ID 9971 (same as Figure 12) under the 10% sensor coverage. Compared to the model with full sensor coverage, the model can still predict the overall trends of the traffic states, but its predictions generally deviate more from the ground truth, as the task has become more challenging. For example in session 52, the model predicts the speed drop later than the full-coverage case, and in session 69 the model has predicted a false congestion around time step 3.
Figure 16: 30-min-ahead prediction for the road segment with ID 9971 under 10% sensor coverage case.
Sensor coverage is a critical factor for prediction performance. Figure 17 shows the 30-min-ahead prediction MAE when the sensor coverage for both modalities varies simultaneously from 1% up to 100%. Firstly, it is straightforward that the higher the sensor coverage, the lower the prediction errors. Still, because of the ability to make use of correlations in space, time and modalities, HiMSNet is able to achieve similar performance for both the regional and segment-level tasks with only 20% sensor coverage. The error metrics of 10% sensor coverage make it a fair trade-off choice, if the coverage is further reduced down to 1% coverage, the performance of the model is close to the Label Average baseline in Table I. This is consistent with the intuition that when the model nearly has no information in the testing phase, the best it can do is to estimate a value based on the training samples, and such predictions (although not always constant) turn out to have nearly constant predictions and similar errors as the constant value used by Label Average.
Figure 17: 30-min-ahead prediction MAE under different sensor coverage
V Conclusion
This paper investigates the problem of multi-source traffic forecasting, i.e., utilizing drone and loop detector data for road segment-level and regional traffic speed prediction. A flexible baseline model HiMSNet is designed to learn the dependencies in temporal, spatial and data modality dimensions. A well-calibrated simulation dataset SimBarca is created for this study, and the experiments demonstrate the effectiveness of drone data in urban traffic forecasting. The high-quality aerial data supports more accurate traffic forecasting compared to using loop detector data. High traffic demand is found to be a challenging factor that causes more dynamic traffic patterns, and thus increases the difficulty of traffic forecasting. Our experiments have shown the importance of spatial correlation and have highlighted the necessity of the drone data modality in traffic forecasting from various perspectives, especially in a real-world scenario with partial and noisy data.
In the future, this work can be improved in several directions. First, more advanced neural network architectures can be explored to better capture the complex dependencies in the traffic data, as the current model is designed to be concise and serve as a baseline. Second, the drone flight in the partial and noisy case can be carefully designed or intelligently optimized to provide best possible observation for traffic forecasting. Third, more data modalities could be included, such as vehicle flow and density. These data are also practically available and can be integrated into the model. We hope our findings can pave the way for more accurate and reliable traffic forecasting in complex urban transportation systems.